# Minimum gas mass accreted by spinning intermediate-mass black holes in stellar clusters

### v1

@@ Abstract
The spin of intermediate-mass black holes (IMBHs) growing through repeated black hole mergers in stellar clusters statistically asymptotes to zero. Putative observations of IMBHs with dimensionless spin parameter $\chi\gtrsim 0.6$ would require a phase of coherent gas accretion to spin up the black hole. We estimate the amount of gas necessary to produce a given IMBH spin. If the observed IMBH mass and spin are $M\gtrsim 1000~{}M_{\odot}$ and $\chi\gtrsim 0.6$, respectively, the IMBH must have coherently accreted at least $\sim 100~{}M_{\odot}$ of gas. In this scenario, as long as the spin is not maximal, the IMBH can only accrete at most half of its mass. Our estimates can constrain the relative contribution of accretion and mergers to the growth of IMBHs in dense stellar environments.
@@ I Introduction
In the supermassive BH regime, spin measurements via X-ray reflection spectroscopy point toward a population of objects with dimensionless spin magnitude $\chi=cJ/(GM^{2})$ (where $M$ and $J$ are the BH mass and angular momentum, respectively) close to the extremal limit $\chi\simeq 1$. However, systematic effects and selection biases for this technique still pose an issue [4]. Gravitational wave (GW) observations provide an alternative method for measuring the spins of merging BHs and extracting their properties. Current observations of merging stellar mass BHs by LIGO-Virgo-KAGRA hint at a population of slowly spinning BHs [5] (but see e.g. Refs. [6, 7] for caveats).
@@ II Black hole mass and spin evolution
II.1 Repeated black-hole mergers
@@
We choose $\alpha=-2$ as a fiducial value, corresponding to equal BH mass per logarithmic mass bin [39], but we will also explore the effect of assuming $\alpha=-1$. We consider a minimum BH mass $m_{\rm min}=5M_{\odot}$ following Ref. [5] (but see e.g. Ref. [40] for uncertainties on this parameter), and a maximum mass $m_{\rm max}=50M_{\odot}$ which is motivated by models of initial BH mass spectra with pair-instability physics in metal-poor environments (see e.g. Ref. [41]). Stellar evolution considerations and GW observations have been used to suggest that BHs may form with spin magnitudes as low as $\sim 10^{-2}$ [42]. We assume all BHs to form with zero natal spin in our fiducial model ($\chi_{0}=0$), but this assumption does not affect the IMBH mass-spin relation, because the memory of the initial spin is quickly erased within a few merger generations. We compute the final mass and final spin using the expressions of Ref. [43], which are based on an interpolation between numerical relativity fits and the test-particle limit. Furthermore, the spin direction of the growing BH is isotropically sampled, modeling the stochasticity in the evolution of the BH spin as its mass grows through successive mergers in a roughly spherically symmetric environment like a cluster (the picture would be quite different in e.g. an axisymmetric environment like a disk [44, 45]).
@@
Figure 1 shows the mass-spin evolution of IMBHs as they grow through repeated mergers. We denote by $M$ and $\chi$ the mass and spin of the growing BH, respectively. We illustrate the fiducial model described above, and then we vary each of the parameters ($\alpha,m_{\rm max},\chi_{0}$) one at a time while keeping the other parameters fixed. As expected, the seed BH spin does not affect the mass-spin relation in the IMBH regime ($M>100M_{\odot}$), as seen by comparing the extreme cases where all seed BHs are nonspinning (black) and maximally spinning (red). The model with $\alpha=-1$ (green) represents a shallower mass spectrum, which produces more heavy BHs (with mass close to $m_{\rm max}$) than the fiducial model. In this scenario, at a fixed mass $M$, the spin has a higher value and takes longer to drop to zero. By a similar argument, when we chose $m_{\rm max}=20M_{\odot}$ (blue), the BH spin magnitude decreases faster. In the scenario where $m_{\rm max}=\infty$, the spin evolves the slowest because there is always a chance for the IMBH to pair up with a BH of similar mass, in which case the remnant spin will be around $\chi\sim 0.7$.
Figure 2: Mass-spin evolutionary history example for an IMBH growing by repeated BH mergers (red line, mass gained $\Delta M_{\rm mer}$) followed by a phase of coherent gas accretion (blue line, mass accreted $\Delta M_{\rm acc}$). The blue-dashed branch corresponds to the case of a retrograde accretion phase that flips the spin direction. The solid red-filled region corresponds to the area that encompasses $99\%$ of realizations of the fiducial hierarchical merger model.
@@ II.2 Coherent gas accretion
The direction of $\chi_{T}$ and $\chi_{f}$ may differ if the IMBH undergoes an initial phase of retrograde accretion that flips the spin followed by a subsequent prograde accretion phase, in which the spin grows again. In principle, it is impossible to know whether the spin of the IMBH flipped in the past just by measuring the spin magnitude $\chi_{f}$. Thus, the two scenarios are degenerate, and in our fiducial forward modeling of BH growth, we give equal probability to both cases once the IMBH reaches the transition point. We also consider a case where we give more weight to the prograde scenario with probability $p_{\rm prog}=0.8$. In practice, this parameter can be inferred from measuring the mass, spin, and relative direction of spin and disk angular momentum for a population of accreting IMBHs. Naively, $p_{\rm prog}=N_{\rm prog}/N$, where $N_{\rm prog}$ is the number of IMBHs observed in their prograde accretion phase, and $N$ is the size of the population. However, depending on the accretion rate, IMBHs may be more likely to be observed during their prograde phase, hence selection effects need to be taken into account when measuring $p_{\rm prog}$ from a population.
@@
Figure 3: Lower bound $\Delta M_{\rm acc}^{\rm min}$ on the gas mass that must be accreted by an IMBH with observed mass $M_{f}$ and spin $\chi_{f}$. The black solid (dashed) line corresponds to the upper boundary of the area that encompasses the $99\%$ ($90\%$) of forward hierarchical merger histories.
@@ II.3 Minimum mass accreted
Implementing the methodology above, we calculate the lower bound of the gas mass $\Delta M_{\rm acc}^{\rm min}$ that must be accreted by a putative IMBH observed in the forbidden region of the fiducial model. Essentially, we integrate Eq. (1) backward with a negative mass step ($dM<0$) from $(M_{f},\chi_{f})$ until $(M,\chi)$ reaches the upper boundary of the area encompassing $99\%$ ($90\%$) of hierarchical merger realizations (cf. Fig. 2). If $M_{T}$ is the mass value of that boundary point, then $\Delta M_{\rm acc}^{\rm min}=|M_{f}-M_{T}|$, and we compute this on a fine grid of final conditions in the $M$$\chi$ plane. We show the results in Fig. 3. Most notably, we find that if $M_{f}\gtrsim 1000M_{\odot}$ and $\chi_{f}\gtrsim 0.6$, the IMBH must coherently accrete at least $\sim 100M_{\odot}$ of gas, while an IMBH with $(M_{f},\chi_{f})\simeq(5000M_{\odot},0.9)$ must have accreted at least $\sim 1500M_{\odot}$ of gas. Combined with precise measurements of IMBH properties, these estimates can be used to constrain the amount of gas in the environment of highly spinning IMBHs. In the next section, we infer the accreted gas mass, including in particular its most likely value and the upper bound on $\Delta M_{\rm acc}$.
@@ III Distribution of accreted gas mass
To infer $\Delta M_{\rm mer}$ and $\Delta M_{\rm acc}$, we use Markov Chain Monte Carlo (MCMC) to explore different forward evolutionary histories that are consistent with the final observed IMBH parameters $(M_{f},\chi_{f})$ accounting for some measurement errors $(\delta M_{f},\delta\chi_{f})$. We describe the computation of our likelihood function in Sec. III.1 below, and in Sec. III.2 we present our inference results.
@@ III.2 MCMC results
Figure 4: Posterior distribution inferred for $M_{f}=5000M_{\odot}$ and $\chi_{f}=0.9$ (left panels), $0.6$ (right panels), assuming fiducial hyperparameters. The relative measurement error on both $M_{f}$ and $\chi_{f}$ is fixed to 1% (top panels) and 10% (bottom panels). The 1-sigma (39%) and 2-sigma (86%) contours are shown. The probability for a purely prograde accretion phase $p_{\rm prog}$ is set to $0.5$; in the top-left panel, we compare this against the case with $p_{\rm prog}=0.8$ (grey curves).
@@
In Fig. 4 we show the posterior distribution for a putative IMBH detection with $M_{f}=5000M_{\odot}$ and either $\chi_{f}=0.9$ (left-column panels) or $\chi_{f}=0.6$ (right-column panels) assuming the fiducial hyperparameters. The top (bottom) row assumes a relative measurement error of $1\%$ ($10\%$) on the IMBH parameters. Furthermore, in the top left panel, we compare $p_{\rm prog}=0.5$ (red) with $p_{\rm prog}=0.8$ (cyan). In Appendix B we discuss the case where $M_{f}=500M_{\odot}$.
@@ IV Conclusions
On the modeling side, the escape velocity of globular clusters is too low to allow for the hierarchical merger products to be retained in the cluster (see e.g. Ref. [18], where the escape velocity needs to exceed $\approx 300\,\rm km\,s^{-1}$ for an IMBH to be produced by hierarchical mergers). Therefore, nuclear star clusters with escape velocity in the hundreds of $\rm km\,s^{-1}$ are more favorable environments for producing IMBHs through repeated mergers [20]. Perturbations on galactic scales may induce gas inflows into the central regions hosting the nuclear star cluster and the IMBH. This may lead to an episode of coherent gas accretion that could spin up the IMBH, as modeled here. If these IMBHs are common in the nuclei of dwarf galaxies, their associated accretion signatures could be identified as low-luminosity active galactic nuclei [55].
@@
Channels other than the scenario examined in this work may also contaminate the high-mass, high-spin region of the IMBH population. A merger of two IMBHs with comparable masses will bring the spin to a high value ($\chi\sim 0.7$) independently of the spin of the IMBH progenitors. Reference [58] shows that the remnant spin distribution for different progenitor spins and mass ratios still displays an exclusion region above $\chi\sim 0.8$, which we also identified in the $m_{\rm max}=\infty$ case in Fig. 1.
@@ Acknowledgements.
K.K. is supported by the Onassis Foundation - Scholarship ID: F ZT 041-1/2023-2024. K.K., L.R. and E.B. are supported by NSF Grants No. AST-2307146, No. PHY-2207502, No. PHY-090003, and No. PHY-20043, by NASA Grants No. 20-LPS20-0011 and No. 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation, and by the Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (rockfish.jhu.edu), which is supported by the NSF Grant No. OAC-1920103. D.G. is supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. Computational work was performed at CINECA with allocations through INFN and Bicocca.
Figure 6: The left panel shows trace plots of a single walker for each run presented in Fig. 4. Each color corresponds to a different case of Fig. 4: red (cyan) to the top left panel with $p_{\rm prog}=0.5$ ($p_{\rm prog}=0.8$), green to the top-right, blue to the bottom left, and yellow to the bottom right. The right panel shows the autocorrelation function of $\Delta M_{\rm acc}$ for all walkers. The dashed magenta lines represent the best-fit exponential model with the integrated autocorrelation time shown in the reported labels.

### v2

@@ Abstract
The spin of intermediate-mass black holes (IMBHs) growing through repeated black hole mergers in stellar clusters statistically asymptotes to zero. Putative observations of IMBHs with dimensionless spin parameter $\chi\gtrsim 0.6$ would require a phase of coherent gas accretion to spin up the black hole. We estimate the amount of gas necessary to produce a given IMBH spin. If the observed IMBH mass and spin are $M\gtrsim 1000M_{\odot}$ and $\chi\gtrsim 0.6$, respectively, the IMBH must have coherently accreted at least $\sim 100M_{\odot}$ of gas. In this scenario, as long as the spin is not maximal, the IMBH can only accrete half of its mass at most. Our estimates can constrain the relative contribution of accretion and mergers to the growth of IMBHs in dense stellar environments.
@@ I Introduction
In the supermassive BH regime, spin measurements via x-ray reflection spectroscopy point toward a population of objects with dimensionless spin magnitude $\chi=cJ/(GM^{2})$ (where $M$ and $J$ are the BH mass and angular momentum, respectively) close to the extremal limit $\chi\simeq 1$. However, systematic effects and selection biases for this technique still pose an issue [4]. Gravitational wave (GW) observations provide an alternative method for measuring the spins of merging BHs and extracting their properties. Current observations of merging stellar-mass BHs by LIGO-Virgo-KAGRA hint at a population of slowly spinning BHs [5] (but see, e.g., Refs. [6, 7] for caveats).
@@ II Black hole mass and spin evolution
II.1 Repeated black hole mergers
@@
We choose $\alpha=-2$ as a fiducial value, corresponding to equal BH mass per logarithmic mass bin [39], but we will also explore the effect of assuming $\alpha=-1$. We consider a minimum BH mass $m_{\rm min}=5M_{\odot}$ following Ref. [5] (but see e.g. Ref. [40] for uncertainties on this parameter), and a maximum mass $m_{\rm max}=50M_{\odot}$ which is motivated by models of initial BH mass spectra with pair-instability physics in metal-poor environments (see e.g. Ref. [41]). Stellar evolution considerations and GW observations have been used to suggest that BHs may form with spin magnitudes as low as $\sim 10^{-2}$ [42]. We assume all BHs to form with zero natal spin in our fiducial model ($\chi_{0}=0$), but this assumption does not affect the IMBH mass-spin relation, because the memory of the initial spin is quickly erased within a few merger generations. We compute the final mass and final spin using the expressions of Ref. [43], which are based on an interpolation between numerical relativity fits and the test-particle limit. Furthermore, the spin direction of the growing BH is isotropically sampled, modeling the stochasticity in the evolution of the BH spin as its mass grows through successive mergers in a roughly spherically symmetric environment like a cluster (the picture would be quite different in, e.g., an axisymmetric environment like a disk [44, 45]).
@@
Figure 1 shows the mass-spin evolution of IMBHs as they grow through repeated mergers. We denote by $M$ and $\chi$ the mass and spin of the growing BH, respectively. We illustrate the fiducial model described above, and then we vary each of the parameters ($\alpha,m_{\rm max},\chi_{0}$) one at a time while keeping the other parameters fixed. As expected, the seed BH spin does not affect the mass-spin relation in the IMBH regime ($M>100M_{\odot}$), as seen by comparing the extreme cases where all seed BHs are nonspinning (black) and maximally spinning (red). The model with $\alpha=-1$ (green) represents a shallower mass spectrum, which produces more heavy BHs (with mass close to $m_{\rm max}$) than the fiducial model. In this scenario, at a fixed mass $M$, the spin has a higher value and takes longer to drop to zero. By a similar argument, when we choose $m_{\rm max}=20M_{\odot}$ (blue), the BH spin magnitude decreases faster. In the scenario where $m_{\rm max}=\infty$, the spin evolves the slowest because there is always a chance for the IMBH to pair up with a BH of similar mass, in which case the remnant spin will be around $\chi\sim 0.7$.
Figure 2: Mass-spin evolutionary history example for an IMBH growing by repeated BH mergers (red line, mass gained $\Delta M_{\rm mer}$) followed by a phase of coherent gas accretion (blue line, mass accreted $\Delta M_{\rm acc}$). The blue dashed branch corresponds to the case of a retrograde accretion phase that flips the spin direction. The solid red-filled region corresponds to the area that encompasses $99\%$ of realizations of the fiducial hierarchical-merger model.
@@ II.2 Coherent gas accretion
The direction of $\chi_{T}$ and $\chi_{f}$ may differ if the IMBH undergoes an initial phase of retrograde accretion that flips the spin followed by a subsequent prograde accretion phase, in which the spin grows again. In principle, it is impossible to know whether the spin of the IMBH flipped in the past just by measuring the spin magnitude $\chi_{f}$. Thus, the two scenarios are degenerate, and in our fiducial forward modeling of BH growth, we give equal probability to both cases once the IMBH reaches the transition point. We also consider a case where we give more weight to the prograde scenario with probability $p_{\rm prog}=0.8$ (while illustrative, this specific choice is arbitrary; we are not aware of a reference or simulation that attempts to estimate the value of $p_{\rm prog}$). In practice, this parameter can be inferred from measuring the mass, spin, and relative direction of spin and disk angular momentum for a population of accreting IMBHs. Naively, $p_{\rm prog}=N_{\rm prog}/N$, where $N_{\rm prog}$ is the number of IMBHs observed in their prograde accretion phase, and $N$ is the size of the population. However, depending on the accretion rate, IMBHs may be more likely to be observed during their prograde phase, hence selection effects need to be taken into account when measuring $p_{\rm prog}$ from a population.
@@
Figure 3: Lower bound $\Delta M_{\rm acc}^{\rm min}$ on the gas mass that must be accreted by an IMBH with observed mass $M_{f}$ and spin $\chi_{f}$. The black solid (dashed) line corresponds to the upper boundary of the area that encompasses the $99\%$ ($90\%$) of forward hierarchical-merger histories.
@@ II.3 Minimum mass accreted
Implementing the methodology above, we calculate the lower bound of the gas mass $\Delta M_{\rm acc}^{\rm min}$ that must be accreted by a putative IMBH observed in the forbidden region of the fiducial model. Essentially, we integrate Eq. (1) backward with a negative mass step ($dM<0$) from $(M_{f},\chi_{f})$ until $(M,\chi)$ reaches the upper boundary of the area encompassing $99\%$ ($90\%$) of hierarchical-merger realizations (cf. Fig. 2). If $M_{T}$ is the mass value of that boundary point, then $\Delta M_{\rm acc}^{\rm min}=|M_{f}-M_{T}|$, and we compute this on a fine grid of final conditions in the $M$$\chi$ plane. We show the results in Fig. 3. Most notably, we find that if $M_{f}\gtrsim 1000M_{\odot}$ and $\chi_{f}\gtrsim 0.6$, the IMBH must coherently accrete at least $\sim 100M_{\odot}$ of gas, while an IMBH with $(M_{f},\chi_{f})\simeq(5000M_{\odot},0.9)$ must have accreted at least $\sim 1500M_{\odot}$ of gas. Combined with precise measurements of IMBH properties, these estimates can be used to constrain the amount of gas in the environment of highly spinning IMBHs. In the next section, we infer the accreted gas mass, including, in particular, its most likely value and the upper bound on $\Delta M_{\rm acc}$.
@@ III Distribution of accreted gas mass
To infer $\Delta M_{\rm mer}$ and $\Delta M_{\rm acc}$, we use the Markov chain Monte Carlo (MCMC) technique to explore different forward evolutionary histories that are consistent with the final observed IMBH parameters $(M_{f},\chi_{f})$ accounting for some measurement errors $(\delta M_{f},\delta\chi_{f})$. We describe the computation of our likelihood function in Sec. III.1 below, and in Sec. III.2 we present our inference results.
@@ III.2 MCMC results
Figure 4: Posterior distribution inferred for $M_{f}=5000M_{\odot}$ and $\chi_{f}=0.9$ (left), $0.6$ (right), assuming fiducial hyperparameters. The relative measurement error on both $M_{f}$ and $\chi_{f}$ is fixed to 1% (top) and 10% (bottom). The 1-$\sigma$ (39%) and 2-$\sigma$ (86%) contours are shown. The probability for a purely prograde accretion phase $p_{\rm prog}$ is set to $0.5$; in the top left panel, we compare this against the case with $p_{\rm prog}=0.8$ (gray curves).
@@
In Fig. 4 we show the posterior distribution for a putative IMBH detection with $M_{f}=5000M_{\odot}$ and either $\chi_{f}=0.9$ (left column panels) or $\chi_{f}=0.6$ (right column panels) assuming the fiducial hyperparameters. The top (bottom) row assumes a relative measurement error of $1\%$ ($10\%$) on the IMBH parameters. Furthermore, in the top left panel, we compare $p_{\rm prog}=0.5$ (red) with $p_{\rm prog}=0.8$ (cyan). In Appendix B we discuss the case where $M_{f}=500M_{\odot}$.
@@ IV Conclusions
On the modeling side, the escape velocity of globular clusters is too low to allow for the hierarchical-merger products to be retained in the cluster (see, e.g., Ref. [18], where the escape velocity needs to exceed $\approx 300\,\rm km\,s^{-1}$ for an IMBH to be produced by hierarchical mergers). Therefore, nuclear star clusters with escape velocity in the hundreds of $\rm km\,s^{-1}$ are more favorable environments for producing IMBHs through repeated mergers [20]. Perturbations on galactic scales may induce gas inflows into the central regions hosting the nuclear star cluster and the IMBH. This may lead to an episode of coherent gas accretion that could spin up the IMBH, as modeled here. If these IMBHs are common in the nuclei of dwarf galaxies, their associated accretion signatures could be identified as low-luminosity active galactic nuclei (AGN) [55].
@@
Our method is purposefully agnostic of any particular astrophysical scenario, although we envision the centers of dwarf galaxies with mini-AGN, such as NGC 4395 [58], as potential environments where this scenario could be realized. A decoupling of the hierarchical BH merging and gas accretion processes could occur if gas inflow episodes drive large amounts of gas in the nuclear star cluster region of the dwarf galaxy at some later stage in the nuclear star cluster’s evolution. In this paper, we have not attempted to carry out detailed simulations of these complex astrophysical processes. We note that, if the BH accretes at the Eddington rate, the accretion timescale is set by the Salpeter time $\tau_{S}\approx 50\,\rm Myr$. If the time between successive mergers, written as the inverse of the merger rate $\Gamma_{m}^{-1}$, is larger than $\tau_{S}$ at the time of the gas inflow episode, then the two processes can be decoupled. To compute the merger rate, we write $\Gamma_{m}\approx n_{\rm BH}\sigma_{m}v$ where $n_{\rm BH}$, $\sigma_{m}$, and $v$ are the number density of stellar-mass BHs, merger cross section, and BH three-dimensional velocity dispersion, respectively. If we use the two-body capture cross section in Eq. (4) of Ref. [59], with $m_{1}\gg m_{2}$ the mass of the growing IMBH and a fixed $m_{2}=10M_{\odot}$, we obtain
$\Gamma_{m}^{-1}$ $\approx 297{\,\rm Myr}$
×$\times\left(n_{\rm BH}\over{10^{5}\,{\rm pc}^{-3}}\right)^{-1}\left({m_{1}\over 10^{3}M_{\odot}}\right)^{-12/7}\left({v\over 10\,\rm km\,s^{-1}}\right)^{-11/7}.$ (5)
Thus, the condition $\Gamma_{m}^{-1}>\tau_{S}$ is satisfied for a measured IMBH mass of $\sim 1000M_{\odot}$ in a typical environment with $v\sim 10\,\rm km\,s^{-1}$, as long as the central BH number density $n_{\rm BH}$ does not significantly exceed $\sim 10^{6}\,\rm pc^{-3}$.
Channels other than the scenario examined in this work may also contaminate the high-mass, high-spin region of the IMBH population. A merger of two IMBHs with comparable masses will bring the spin to a high value ($\chi\sim 0.7$) independent of the spin of the IMBH progenitors. Reference [60] shows that the remnant spin distribution for different progenitor spins and mass ratios still displays an exclusion region above $\chi\sim 0.8$, which we also identified in the $m_{\rm max}=\infty$ case in Fig. 1.
@@
Figure 6: Left: trace plots of a single walker for each run presented in Fig. 4. Each color corresponds to a different case of Fig. 4: red (cyan) to the top left panel with $p_{\rm prog}=0.5$ ($p_{\rm prog}=0.8$), green to the top right, blue to the bottom left, and yellow to the bottom right. Right: the autocorrelation function of $\Delta M_{\rm acc}$ for all walkers. The dashed magenta lines represent the best-fit exponential model with the integrated autocorrelation time shown in the reported labels.
@@ Acknowledgements.
K.K. is supported by the Onassis Foundation - Scholarship ID: F ZT 041-1/2023-2024. K.K., L.R. and E.B. are supported by NSF Grants No. AST-2307146, No. PHY-2207502, No. PHY-090003, and No. PHY-20043, by NASA Grants No. 20-LPS20-0011 and No. 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation, and by the Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility [72], which is supported by the NSF Grant No. OAC-1920103. D.G. is supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. Computational work was performed at CINECA with allocations through INFN and Bicocca.
Data availability
No data were created or analyzed in this study.